### v1

# Stability of spherical thin-shell wormholes in scalar-tensor theories

Abstract
In this article, we construct a family of spherically symmetric thin-shell wormholes within scalar-tensor theories of gravity. In the case of wormholes symmetric across the throat, we study the matter content and analyze the stability of the static configurations under radial perturbations. We apply the formalism to two particular examples involving Einstein-Maxwell gravity coupled to a conformal scalar field. In both scenarios, we show that stable configurations are possible for suitable values of the parameters involved.
1 Introduction
General Relativity (GR) is a very successful theory of gravity, which has provided explanations for a variety of astrophysical and cosmological phenomena, and has passed many experimental tests [1, 2, 3]. However, there are some unresolved issues, such as the elusive nature of the dark matter [4, 5] necessary to model various observations, the cosmological data indicating that the late-time expansion of the universe is accelerated [6, 7], which requires the presence of dark energy [5], and the so-called Hubble tension [8, 9]. All of these suggest that GR might not be the final answer. In this context, several theories of modified gravity have appeared in the literature. Among them, the scalar-tensor theories are interesting alternatives to GR, in which gravity is mediated by a tensor and a scalar field [10]. They are generalizations of the Brans-Dicke theory [11]. Several topics have been studied within scalar-tensor gravity, such as gravitational waves [12], the post-Newtonian expansion [13], black holes [14, 15, 16, 17], and cosmological aspects [18]. It is worth mentioning the existence of an ambiguity in scalar-tensor theories related to the distinction between the Einstein and the Jordan frames. There is a long debate about the most suitable physical frame and the (non)equivalence of them [19, 20]. Another proposal is $F(R)$ gravity [21, 22], in which the Lagrangian density is a function of the Ricci scalar $R$. More alternative theories to GR can be found in [22].
In the context of the different theories of gravity, there are solutions of the field equations that represent compact objects, such as black holes and wormholes. The existence of wormholes is still hypothetical; they were first introduced by Einstein and Rosen [23] and studied more systematically since the work by Morris and Thorne [24]. These geometries represent bridges that connect two different regions of the Universe or two different universes across a throat [25], where the flare-out condition is fulfilled. A common issue encountered in traversable wormholes is the presence of the so-called exotic matter, i.e. matter that does not satisfy the energy conditions [25], which is not a physically desirable feature. In GR, the exotic matter is always required near the throat [26, 27, 28]; however, its amount can be reduced [29], but at the cost of increasing tensions in it [30]. Many papers concerning wormholes with scalar fields have appeared in the literature, here we can only mention a few [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. On the other hand, a new spacetime can be obtained by the cut and paste of two geometries, which are joined on their boundaries by applying the suitable junction conditions, firstly found in GR [42, 43]. This method is used to construct the denominated thin-shell wormholes [25], that allows to localize and minimize the amount of exotic matter. With this procedure, the new manifold typically has an infinitesimally thin layer of matter at the joining hypersurface, where the throat is placed. The construction, the stability, and other related topics on thin-shell wormholes have been studied within GR [44, 45, 46, 47, 48, 49, 50], in dilaton gravity [51], in Einstein-Gauss-Bonnet [52, 53], in Brans-Dicke [54], in $f(R)$ theory [55, 56, 57, 58, 59, 60], and also in Palatini $f(R)$ gravity [61], among others.
In this paper, we construct thin-shell wormholes with spherical symmetry in Einstein theory non-minimally coupled to a scalar field. In order to have a proper matching at the shell, we apply the junction conditions corresponding to this theory [62]. For a family of wormholes symmetric across the throat, which is located at the shell, we study their dynamical stability and the presence of exotic matter. In particular, for the construction we adopt two black hole solutions within Einstein-Maxwell gravity coupled to a conformally invariant scalar field, one quadratic with a constant scalar field [17] and the other of fourth order with variable scalar field [14]. The structure of this paper is as follows. In Sect. 2, we present the general construction of the wormholes in scalar-tensor gravity. In Sect. 3, we develop the formalism for the stability of wormholes symmetric across the throat. In Sect. 4, we consider two particular examples of spacetimes with a radial electric field, in which we study the stability and the energy conditions. Finally, we summarize our results in Sect. 5. We use units such that $c=1$.
2 Wormhole construction
In this section, we construct a class of spherically symmetric thin-shell wormholes in scalar-tensor gravity by adopting the so-called junction formalism. We are interested in a class of scalar-tensor theories with a non-minimally coupled real scalar field $\phi$ and a self-interaction potential $U(\phi)$, described by the action in the Jordan frame
$S=\int d^{4}x\,\sqrt{-g}\left(f(\phi)R-\frac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi-U(\phi)\right)+\int d^{4}x\,\sqrt{-g}\mathcal{L}_{M},$ (1)
from which the following field equations are obtained
$2f(\phi)G_{\mu\nu}+g_{\mu\nu}\left(\frac{1}{2}(\nabla\phi)^{2}+U(\phi)\right)-\nabla_{\mu}\phi\nabla_{\nu}\phi-2\nabla_{\mu}\nabla_{\nu}f(\phi)+2g_{\mu\nu}\Box f(\phi)=T_{\mu\nu},$ (2)
$\Box\phi+f^{\prime}(\phi)R-U^{\prime}(\phi)=0,$ (3)
where $T_{\mu\nu}$ is the energy-momentum tensor associated with the other matter field Lagrangian $\mathcal{L}_{M}$.
The junction formalism for these theories, introduced in Ref. [62], establishes the conditions for proper matching at a hypersurface $\Sigma$. We denote the jump of any quantity $\Upsilon$ across $\Sigma$ by
$\Upsilon]\equiv(\Upsilon^{2}-\Upsilon^{1})|_{\Sigma}.$ (4)
The first set of conditions arises from the continuity of the metric and the scalar field across the matching hypersurface, respectively
$g_{\mu\nu}]=0,$ (5)
$\phi]=0,$ (6)
while the second set comes from the Einstein equations and the equation of motion of the scalar field, respectively
$2\varepsilon f(\phi)([K_{\mu\nu}]-h_{\mu\nu}[K])+2Mf^{\prime}(\phi)h_{\mu\nu}=S_{\mu\nu},$ (7)
$M=2\varepsilon f^{\prime}(\phi)[K].$ (8)
In these equations, the prime in $f(\phi)$ represents derivative with respect to $\phi$, $K_{\mu\nu}$ is the second fundamental form, $S_{\mu\nu}$ is the energy-momentum tensor at $\Sigma$, and $M$ is defined by
$M=\varepsilon n^{\mu}[\partial_{\mu}\phi],$ (9)
where $n^{1,2}_{\gamma}$ is the unit normal to $\Sigma$, satisfying $n^{\gamma}n_{\gamma}=\varepsilon$. The hypersurface $\Sigma$ in our case will be timelike and so we set $\varepsilon=1$ from now on (the value $\varepsilon=-1$ corresponds to a spacelike hypersurface). One speaks of a boundary hypersurface when $S_{\mu\nu}\equiv 0$, otherwise of a thin shell of matter.
In our construction, we start from two spherically symmetric spacetimes described by the line elements
$ds_{1,2}^{2}=-A_{1,2}(r)dt_{1,2}^{2}+A_{1,2}(r)^{-1}dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\varphi^{2}),$ (10)
where $r\geq 0$ is the radial coordinate, while $0\leq\theta\leq\pi$ and $0\leq\varphi<2\pi$ are the angular coordinates. We select a radius $a>0$ and remove the region with $r<a$ from each geometry, in order to define the manifolds
$\mathcal{M}^{1,2}=\{X_{1,2}^{\alpha}=(t_{1,2},r,\theta,\varphi)/r\geq a\},$ (11)
and we join them at the surface
$\Sigma\equiv\Sigma^{1,2}=\{X_{1,2}/\mathcal{G}(r)=r-a=0\}.$ (12)
This cut-and-paste procedure has as a result a new geodesically complete manifold $\mathcal{M}=\mathcal{M}^{1}\cup\mathcal{M}^{2}$. The area $4\pi r^{2}$ is minimal at $r=a$, so the flare-out condition is satisfied there and $\mathcal{M}$ describes a wormhole with two regions connected by the throat $\Sigma$ with radius $a$. We can use the proper radial distance $\ell=\pm\int_{a}^{r}\sqrt{1/A_{1,2}(r)}dr$ in order to introduce a global radial coordinate in $\mathcal{M}$, where the (-) and (++) sign corresponds to $\mathcal{M}^{1}$ and to $\mathcal{M}^{2}$ respectively, with the wormhole throat placed at $\ell=0$.
Apart from the coordinates that describe the 4-dimensional spacetime, we introduce the coordinates $\xi^{i}=(\tau,\theta,\varphi)$ on the surface $\Sigma$, with $\tau$ the proper time on the shell. We let the throat radius $a$ to be a function of $\tau$ in order to study of the stability of the shell under radial perturbations. The line element should be continuous across $\Sigma$ as the time coordinates in each side satisfy that $d\tau^{2}=A_{1}(a)^{2}\left(A_{1}(a)+\dot{a}^{2}\right)^{-1}dt_{1}^{2}=A_{2}(a)^{2}\left(A_{2}(a)+\dot{a}^{2}\right)^{-1}dt_{2}^{2}$, where the dot represents the derivative with respect to $\tau$. The induced metric on $\Sigma$ then reads
$ds_{\Sigma}^{2}=-d\tau^{2}+a^{2}(\tau)(d\theta^{2}+\sin^{2}\theta d\varphi^{2}).$ (13)
The first fundamental form $h_{ij}$ or the metric on the shell associated with the two sides of it is given by
$h^{1,2}_{ij}=\left.g^{1,2}_{\mu\nu}\frac{\partial X^{\mu}_{1,2}}{\partial\xi^{i}}\frac{\partial X^{\nu}_{1,2}}{\partial\xi^{j}}\right|_{\Sigma},$ (14)
and the second fundamental form or the extrinsic curvature of the shell $K_{ij}$ is calculated from
$K_{ij}^{1,2}=-n_{\gamma}^{1,2}\left.\left(\frac{\partial^{2}X^{\gamma}_{1,2}}{\partial\xi^{i}\partial\xi^{j}}+\Gamma_{\alpha\beta}^{\gamma}\frac{\partial X^{\alpha}_{1,2}}{\partial\xi^{i}}\frac{\partial X^{\beta}_{1,2}}{\partial\xi^{j}}\right)\right|_{\Sigma},$ (15)
where
$n_{\gamma}^{1,2}=\pm\left\{\left.\left|g^{\alpha\beta}_{1,2}\frac{\partial\mathcal{G}}{\partial X^{\alpha}_{1,2}}\frac{\partial\mathcal{G}}{\partial X^{\beta}_{1,2}}\right|^{-1/2}\frac{\partial\mathcal{G}}{\partial X^{\gamma}_{1,2}}\right\}\right|_{\Sigma},$ (16)
is the unit normal($n^{\gamma}n_{\gamma}=1$) for each side of the shell. For the metrics given by Eq. (10), they read
$n_{\gamma}^{1,2}=\pm\left(-\dot{a},\frac{\sqrt{A_{1,2}(a)+\dot{a}^{2}}}{A_{1,2}(a)},0,0\right).$ (17)
On $\Sigma$, we can use of the orthonormal basis $\{e_{\hat{\tau}}=e_{\tau},e_{\hat{\theta}}=a^{-1}e_{\theta},e_{\hat{\varphi}}=(a\sin\theta)^{-1}e_{\varphi}\}$ for which the first fundamental form is $h^{1,2}_{\hat{\imath}\hat{\jmath}}=\mathrm{diag}(-1,1,1)$, and the components of the second fundamental form are
$K_{\hat{\tau}\hat{\tau}}^{1,2}=\mp\left(\frac{\ddot{a}}{\sqrt{A_{1,2}(a)+\dot{a}^{2}}}+\frac{1}{2}\frac{A^{\prime}_{1,2}(a)}{\sqrt{A_{1,2}(a)+\dot{a}^{2}}}\right)$ (18)
and
$K_{\hat{\theta}\hat{\theta}}^{1,2}=K_{\hat{\varphi}\hat{\varphi}}^{1,2}=\pm\frac{1}{a}\sqrt{A_{1,2}(a)+\dot{a}^{2}},$ (19)
where the prime represents the derivative with respect to $r$. In this basis, we see that the angular components $K_{\hat{\theta}\hat{\theta}}$ and $K_{\hat{\phi}\hat{\phi}}$ are the same and the junction condition (5), i.e. $h_{\hat{\imath}\hat{\jmath}}]=0$, is immediately fulfilled. The jump of the trace of the extrinsic curvature is given by
$K]=\frac{1}{2}\left(\frac{4\left(\sqrt{A_{1}(r)+\dot{a}^{2}}+\sqrt{A_{2}(r)+\dot{a}^{2}}\right)}{a}+\frac{A_{1}^{\prime}(r)+2\ddot{a}}{\sqrt{A_{1}(r)+\dot{a}^{2}}}+\frac{A_{2}^{\prime}(r)+2\ddot{a}}{\sqrt{A_{2}(r)+\dot{a}^{2}}}\right).$ (20)
By using the definition of $M$ introduced in Eq. (9) and considering that the angular components of the normal vector vanish, we obtain the expression
$M=\frac{\dot{a}^{2}\phi^{\prime}_{2}}{\sqrt{A_{2}+\dot{a}^{2}}}-\frac{\dot{a}^{2}\phi^{\prime}_{1}}{\sqrt{A_{1}+\dot{a}^{2}}}+\sqrt{A_{2}+\dot{a}^{2}}\phi^{\prime}_{2}-\sqrt{A_{1}+\dot{a}^{2}}\phi^{\prime}_{1}.$ (21)
The junction condition (8) takes the form
$M+4f^{\prime}(\phi)\left(\frac{\sqrt{A_{1}(r)+\dot{a}^{2}}+\sqrt{A_{2}(r)+\dot{a}^{2}}}{a}+\frac{A_{1}^{\prime}(r)+2\ddot{a}}{4\sqrt{A_{1}(r)+\dot{a}^{2}}}+\frac{A_{2}^{\prime}(r)+2\ddot{a}}{4\sqrt{A_{2}(r)+\dot{a}^{2}}}\right)=0,$ (22)
[…]
$\sigma=-2Mf^{\prime}(\phi)-\frac{4f(\phi)}{a}(\sqrt{A_{1}(r)+\dot{a}^{2}}+\sqrt{A_{2}(r)+\dot{a}^{2}})$ (23)
and
$p=2Mf^{\prime}(\phi)+f(\phi)\left(\frac{2(\sqrt{A_{1}(r)+\dot{a}^{2}}+\sqrt{A_{2}(r)+\dot{a}^{2}})}{a}+\frac{A_{1}^{\prime}(r)+2\ddot{a}}{\sqrt{A_{1}(r)+\dot{a}^{2}}}+\frac{A_{2}^{\prime}(r)+2\ddot{a}}{\sqrt{A_{2}(r)+\dot{a}^{2}}}\right).$ (24)
where $M$ shown in Eq. (21). All the equations above are greatly simplified if the wormhole spacetime is symmetric across the throat, i.e. when the solutions adopted at both sides of the shell are the same.
3 Stability of wormholes symmetric across the throat
For studying the stability of the static configurations, with a throat radius denoted by $a_{0}$, one usually aims to obtain an expression for $\dot{a}^{2}$ and express it in a form of a master equation
$\dot{a}^{2}=-V(a),$ (25)
where $V(a)$ can be understood as a potential. The stable solutions under radial perturbations are those that fulfill the condition $V^{\prime\prime}(a_{0})>0$. In our case, it appears quite cumbersome to solve the dynamics in terms of $\dot{a}^{2}$ in a generalized way. However, our study simplifies for thin-shell wormholes symmetric across the throat. In this case, the two spacetimes that are joined at $\Sigma$ are the same, thus $A_{1}(r)=A_{2}(r)=A(r)$ and $\phi_{1}(r)=\phi_{2}(r)=\phi(r)$. This also implies that $\partial_{\mu}\phi]=0$ and consequently $M=0$ because of Eq. (9). The junction condition (8) then reduces to $[K]=0$, which using Eq. (20) takes the form
$8f^{\prime}(\phi)\left(\frac{\sqrt{A(r)+\dot{a}^{2}}}{a}+\frac{A^{\prime}(r)+2\ddot{a}}{4\sqrt{A(r)+\dot{a}^{2}}}\right)=0.$ (26)
Since $f^{\prime}(\phi)\neq 0$ beyond GR, we can rewrite it as
$4(A(r)+\dot{a}^{2})+a(A^{\prime}(r)+2\ddot{a})=0.$ (27)
By using (27) and that $2\ddot{a}=d\dot{a}^{2}/da$, we can find an expression for $\dot{a}^{2}$ in the form of Eq. (25), in which we understand the following expression as the potential
$V(a)=A(a)-\frac{a_{0}^{4}}{a^{4}}A(a_{0}).$ (28)
The second order Taylor expansion of this potential reads
$V(a)=V(a_{0})+V^{\prime}(a_{0})(a-a_{0})+\frac{V^{\prime\prime}(a_{0})}{2}(a-a_{0})^{2}+\mathcal{O}(a-a_{0})^{3}.$ (29)
It is not difficult to verify that $V(a_{0})=0$ and with the aid of Eq. (27) also that $V^{\prime}(a_{0})=0$, while the second derivative of the potential evaluated at $a_{0}$ takes the form
$V^{\prime\prime}(a_{0})=-20\frac{A(a_{0})}{a_{0}^{2}}+A^{\prime\prime}(a_{0}).$ (30)
The throat radius $a_{0}$ should satisfy the junction condition (27), which reduces to
$4A(a_{0})+a_{0}A^{\prime}(a_{0})=0.$ (31)
The perfect fluid at the shell has an energy density (23) and a pressure (24) whose expressions in the static case read, respectively
$\sigma_{0}=-\frac{8f(\phi_{0})}{a_{0}}\sqrt{A(a_{0})}$ (32)
and
$p_{0}=2f(\phi_{0})\left(\frac{2\sqrt{A(a_{0})}}{a_{0}}+\frac{A^{\prime}(a_{0})}{\sqrt{A(a_{0})}}\right),$ (33)
where $\phi_{0}=\phi(a_{0})$. By using Eq. (31) we can easily see that
$p_{0}=-\frac{4f(\phi_{0})}{a_{0}}\sqrt{A(a_{0})},$ (34)
so the equation of state is fixed, having the form $p_{0}=\sigma_{0}/2$. At this point, it is worthy to emphasize that the relation between the jump of the trace of the second fundamental form and the scalar field given by Eq. (8), which for this class of wormholes results in $[K]=0$, leads to the constraint on the shell radius $a_{0}$ shown in Eq. (31). This constraint, not present in GR, also fixes the equation of state at $\Sigma$ in the presence of the scalar field.
Another important aspect to analyze is the type of matter that constitutes the wormhole, which can be studied in terms of the weak energy condition (WEC). This condition, given by the inequalities
$\sigma_{0}\geq 0\qquad\mathrm{and}\qquad\sigma_{0}+p_{0}\geq 0,$ (35)
determines when the matter at the throat satisfies WEC (dubbed normal) or not (exotic). The non-minimal coupling function $f(\phi)$ has to be positive everywhere in order that the effective gravitational constant satisfy $G_{eff}\propto 1/f(\phi_{0})>0$ and (from the quantum point of view) avoid the graviton to be a ghost [63]. Considering Eq. (32) and adopting $f(\phi_{0})>0$, we realize that the energy density at the throat is always a negative quantity and this implies the existence of exotic matter for any choice of the parameters. This fact has actually been proved for any symmetric thin-shell wormhole configuration in any scalar-tensor theory [63]. In other terms, the price to pay for having matter satisfying WEC at the throat, which requires $f(\phi_{0})<0$, is to have $G_{eff}<0$ and a ghost graviton.
4 Wormholes with a radial electric field
In this section, we consider two examples of interest. They correspond to particular cases of Einstein-Maxwell theories coupled to a scalar field. We adopt a solution from each theory to construct the corresponding thin-shell wormholes and we analyze the stability of the static configurations and the fulfillment of the WEC.
4.1 Constant scalar field
The action for Einstein-Maxwell theory with a conformally coupled self interacting scalar field $\phi$ introduced in [17] is given by Eq. (1) with
$f(\phi)=\frac{1}{16\pi G}-\frac{\phi^{2}}{12},$ (36)
$U(\phi)=0,$ (37)
$\mathcal{L}_{M}=-\frac{1}{16\pi}F_{\mu\nu}F^{\mu\nu}.$ (38)
where $F_{\mu\nu}=\partial_{\mu}\mathcal{A}_{\nu}-\partial_{\nu}\mathcal{A}_{\mu}$ is the electromagnetic tensor in terms of the vector potential $\mathcal{A}_{\mu}$. The corresponding gravitational, scalar, and electromagnetic field equations are respectively
$G_{\mu\nu}=8\pi G\left(T^{S}_{\mu\nu}+T^{EM}_{\mu\nu}\right),$ (39)
$\Box\phi=\frac{1}{6}R\phi,$ (40)
$\partial_{\mu}(\sqrt{-g}F^{\mu\nu})=0,$ (41)
where the sum of
$T^{S}_{\mu\nu}=\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{2}g_{\mu\nu}\partial_{\sigma}\phi\partial^{\sigma}\phi+\frac{1}{6}(g_{\mu\nu}\Box-\nabla_{\mu}\nabla_{\nu}+G_{\mu\nu})\phi^{2},$ (42)
and
$T^{EM}_{\mu\nu}=\frac{1}{\mu_{0}}(F_{\mu\rho}F_{\nu}^{\;\rho}-\frac{1}{4}g_{\mu\nu}F_{\rho\sigma}F^{\rho\sigma})$ (43)
is the total energy-momentum tensor $T_{\mu\nu}$.
For the construction of the thin-shell wormholes, we are interested in the static hairy Reissner-Nordström black hole, which is a solution of the field equations with a conformally coupled constant scalar field. In this case, adopting units so that $G=1$ and $\mu_{0}/(4\pi)=1$, the only non-zero component of the electromagnetic field that solves (41) can be expressed as
$\partial_{r}A_{t}=F_{rt}=\frac{Q}{r^{2}},$ (44)
and the metric and scalar field which are solutions of (39) and (40), respectively read [17]
$ds^{2}=-\left(1-\frac{2m}{r}+\frac{Q^{2}+s}{r^{2}}\right)dt^{2}+\left(1-\frac{2m}{r}+\frac{Q^{2}+s}{r^{2}}\right)^{-1}dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})\,$ (45)
$\phi=\pm\sqrt{\frac{6}{8\pi}}\sqrt{\frac{s}{s+Q^{2}}},$ (46)
where $m$ is the mass, $Q$ is the electric charge, and $s$ the scalar hair. The sign of $Q$ does not affect the metric nor the scalar field (it only has influence on the direction of the radial electric field), so we adopt $Q\geq 0$ without losing generality. Because of Eq. (46), for positive values of $s$, we see that $Q$ can take any value, while for negative values of $s$, we should have $Q^{2}<-s$. For $s=0$, there is a null scalar field and we recover the Reissner-Nordström solution of GR, in which $Q$ has no restrictions. The horizons are given by
$r_{\pm}=m\pm\sqrt{m^{2}-Q^{2}-s},$ (47)
when $m^{2}-Q^{2}-s\geq 0$. The plus sign corresponds to the event horizon $r_{h}$ and the minus sign to a Cauchy horizon. For $m^{2}-Q^{2}-s<0$ no horizons are present and there is a naked singularity at the origin, because the Kretschmann invariant $\mathcal{K}=R_{\alpha\beta\gamma\delta}R^{\alpha\beta\gamma\delta}$ (quadratic in the Riemann tensor)
$\mathcal{K}=\frac{48(Q^{2}-mr+s)^{2}}{r^{8}}$ (48)
diverges. When $m^{2}-Q^{2}-s=0$ the two horizons coincide (extreme case). Note that a necessary condition for the presence of the horizons is that $m^{2}\geq s$; in this case there is a critical value of the charge $Q_{c}$, which reads
$Q_{c}=\sqrt{m^{2}-s}$ (49)
so that there are two horizons for $Q<Q_{c}$, one horizon when $Q=Q_{c}$, and no horizons if $Q>Q_{c}$.
Figure 1: Stability of a spherical thin-shell wormholes symmetric across the throat, in Einstein-Maxwell theory with a conformally coupled and constant scalar field, for different values of the hair parameter $s/m^{2}$, with $m$ the mass. The solid line represents the stable static configurations with radius $a_{0}/m$ and charge $Q/m$, while the dashed line shows the unstable solutions. The dark gray regions have no physical meaning (see text) and in the light gray ones the scalar field is not real.
In order to obtain a traversable wormhole, the throat radius should satisfy $a_{0}>r_{h}$ when $Q\leq Q_{c}$, which removes the horizon and the singularity at the origin, and $a_{0}>0$ if $Q>Q_{c}$, which takes away the naked singularity at the origin. We use the formalism of the previous section to analyze the behavior of the static configurations. The possible values of $a_{0}$ are those which satisfy Eq. (31) and the stability is determined by Eq. (30). The scalar field shown in Eq. (46) should be real, which restricts the value of charge to $Q^{2}<-s$ when $s<0$.
We present our results graphically. In Fig. 1, we study the existence and stability of thin-shell wormholes by displaying $a_{0}/m$ against $Q/m$ for different values of the parameter $s/m^{2}$. The black curve corresponds to the solutions that satisfy (31), where the stable configurations are drawn with a solid line and the unstable ones with a dashed line. The dark gray areas show the regions without physical meaning which are removed, while the light gray ones are those where the scalar field is not real. In more detail, for negative $s/m^{2}$, represented by the plot with $s/m^{2}=-1$, we find two solutions for $Q>Q_{c}$, the larger one is stable, while the other is unstable, but in both the scalar field is not real so they have to be discarded. For $0<s/m^{2}<1$, represented by $s/m^{2}=0.5$, two solutions are found when $Q>Q_{c}$, a stable, larger one and an unstable, smaller one. The region without physical meaning shrinks to finally disappear as we increase the value of $s/m^{2}$. For $s/m^{2}=1$, there are two solutions for small values of $Q/m^{2}$; the larger one is stable while the other is unstable. In this case, there are no horizons in the original metric for $Q\neq 0$, and for the case of $Q=0$, there is a horizon where the solutions fall on its border or into it. For $1<s/m^{2}\leq 9/8$, represented by $s/m^{2}=1.1$ the behavior is similar to the case with $s/m^{2}=1$. Finally, if $s/m^{2}\geq 9/8$, one can see that Eq. (31) has no real solutions, so our construction is not possible.
Let us analyze the matter content at the throat. When $f(\phi_{0})>0$, that is $s/(s+Q^{2})<1$, the WEC is not fulfilled since $\sigma_{0}<0$ in Eq. (32). If $f(\phi_{0})\leq 0$, that is $s/(s+Q^{2})\geq 1$, the WEC is satisfied, because $\sigma_{0}\geq 0$ and $\sigma_{0}+p_{0}\geq 0$, with $p_{0}$ given by Eq. (34). When $s>0$ we can see that the matter is always exotic, while for $s<0$, if $Q^{2}<-s$ it is normal and if $s<Q^{2}$ is exotic but with a non-real scalar field. Outside the shell, i.e. in the bulk, the total energy-momentum in an orthonormal frame reads [17]
$T^{\hat{\mu}\hat{\nu}}=\frac{Q^{2}+s}{r^{4}}\mathrm{diag}\left(1,-1,1,1\right),$ (50)
which for a perfect fluid takes the form $T^{\hat{\mu}\hat{\nu}}=\mathrm{diag}(\rho,p_{1},p_{2},p_{3})$, with $\rho$ the volume energy density and $p_{i}$ the pressures. It is easy to check that $\rho\geq 0$ and $\rho+p_{i}\geq 0$ if $Q^{2}+s\geq 0$, so the WEC is satisfied at the bulk in this case. In particular, when $s>0$ the matter is always normal, while for $s<0$ if $Q^{2}\leq-s$ it is exotic and if $s\leq Q^{2}$ is normal but with a non-real scalar field.
In brief, we have found that the configurations with $s<0$ are not physically relevant and those with $s>0$ present interesting results. In particular, if $0<s/m^{2}<1$ there is a pair of solutions with large values of $Q/m$, i.e. $Q>Q_{c}$, one stable and the other unstable, while for $1\leq s<9/8$ the stable and the unstable configurations correspond to small values of $Q/m$. In both scenarios, the WEC is satisfied at the bulk, but not at the throat.
4.2 Variable scalar field
A more general action of the form (1) for Einstein-Maxwell theory with a conformally coupled self interacting scalar field $\phi$ has been presented in [14], with $f(\phi)$ given again by Eq. (36), the electromagnetic Lagrangian by Eq. (38), and
$U(\phi)=\frac{\Lambda}{16\pi G}+\alpha\phi^{4},$ (51)
where $\Lambda$ is the cosmological constant and $\alpha$ is a constant. We notice that this action involves a term of higher order in the scalar field in addition to the one that appears the previous example. The field equations are
$G_{\mu\nu}+\Lambda g_{\mu\nu}=8\pi G\left(T_{\mu\nu}^{\phi}+T_{\mu\nu}^{EM}\right),$ (52)
$\Box\phi-\frac{1}{6}R\phi-4\alpha\phi^{3}=0,$ (53)
and Eq. (41) for the electromagnetic field. The total energy-momentum tensor $T_{\mu\nu}$ is the sum of
$T_{\mu\nu}^{\phi}=\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{2}g_{\mu\nu}g^{\alpha\beta}\partial_{\alpha}\phi\partial_{\beta}\phi+\frac{1}{6}\left[g_{\mu\nu}\Box-\nabla_{\mu}\nabla_{\nu}+G_{\mu\nu}\right]\phi^{2}-\alpha g_{\mu\nu}\phi^{4}$ (54)
and $T_{\mu\nu}^{EM}$ shown in Eq. (43).
We are interested again in a charged spacetime whose metric is solution of Eq. (52), having the metric [14]
$ds^{2}=-\left[-\frac{\Lambda}{3}r^{2}+\left(1-\frac{m}{r}\right)^{2}\right]dt^{2}+\left[-\frac{\Lambda}{3}r^{2}+\left(1-\frac{m}{r}\right)^{2}\right]^{-1}dr^{2}+r^{2}d\Omega^{2}\,$, (55)
where $m$ is the mass. We have again used units that $G=1$ and $\mu_{0}/(4\pi)=1$. It is peculiar that the charge does not appear explicitly in the metric, in contrast to the standard Reissner-Nordström spacetime. Instead, the charge has a fixed value determined by the other parameters. Particularly, the charge-to-mass ratio is
$\left(\frac{Q}{m}\right)^{2}=1+\frac{2\pi\Lambda}{9\alpha},$ (56)
which imposes the following constraint
$\frac{\Lambda}{\alpha}\geq-\frac{9}{2\pi}.$ (57)
Therefore, the quotient $\Lambda/\alpha$ has a lower bound. The scalar field, which is a solution of Eq. (53), takes the following form
$\phi(r)=\sqrt{\frac{3}{4\pi}}\frac{\sqrt{m^{2}-Q^{2}}}{r-m},$ (58)
or, by using Eq. (56),
$\phi(r)=\sqrt{-\frac{\Lambda}{6\alpha}}\;\frac{m}{r-m},$ (59)
which for being real requires that
$\frac{Q^{2}}{m^{2}}\leq 1,\qquad\frac{\Lambda}{\alpha}\leq 0.$ (60)
The only non-vanishing component of the electromagnetic field is given by (44). Considering that $Q/m$ gets a fixed value for a given $\Lambda/\alpha$, it is useful to introduce a new dimensionless parameter $\epsilon=\Lambda/\alpha$. Thus, the previous expressions can be rewritten in the form
$\left(\frac{Q}{m}\right)^{2}=\left(1+\frac{2\pi\epsilon}{9}\right),$ (61)
$\phi(r)=\sqrt{-\frac{\epsilon}{6}}\frac{m}{r-m},$ (62)
with the constraint
$\frac{9}{2\pi}\leq\epsilon\leq 0.$ (63)
It becomes clear that as long as $\Lambda$ and $\alpha$ have opposite signs, we do not have to worry about their relative behavior when we use the parameter $\epsilon$. Then, for our calculations, we can write the final expressions in terms of $\Lambda$ and $\epsilon$. The information about $\alpha$ is not lost since each value of $\Lambda$ corresponds to a specific value of $\alpha$ through the value of $\epsilon$.
We can now proceed to analyze the solution that we will use for constructing the thin-shell wormholes. This solution presents two possible cases: one with $\Lambda\geq 0$, which describes a static and spherically symmetric black hole, and the other with $\Lambda<0$, which corresponds to a naked singularity. For $\Lambda>0$, there are an inner (Cauchy), an event, and a cosmological horizons, given by
$r_{i}=\frac{\sqrt{3}}{2\sqrt{\Lambda}}\left(-1+\sqrt{1+\frac{4m\sqrt{\Lambda},}{\sqrt{3}}}\right),$ (64)
$r_{h}=\frac{\sqrt{3}}{2\sqrt{\Lambda}}\left(1-\sqrt{1-\frac{4m\sqrt{\Lambda},}{\sqrt{3}}}\right),$ (65)
and
$r_{c}=\frac{\sqrt{3}}{2\sqrt{\Lambda}}\left(1+\sqrt{1-\frac{4m\sqrt{\Lambda}}{\sqrt{3}}}\right),$ (66)
respectively, which satisfy $r_{i}\leq r_{h}\leq r_{c}$. By computing the Kretschmann invariant
$\mathcal{K}=\frac{48m^{2}(r-m)^{2}}{r^{8}},$ (67)
we conclude that the solution presents a singularity at the origin. The horizons $r_{h}$ and $r_{c}$ both exist when $0<\Lambda m^{2}\leq 3/16$, they coincide if $\Lambda m^{2}=3/16$, while for $\Lambda m^{2}>3/16$ they vanish but the inner horizon $r_{i}$ remains. When $\Lambda<0$ the metric function is always positive and thus, there are no horizons and the singularity is naked.
Figure 2: Stability of a spherical thin-shell wormholes symmetric across the throat, in Einstein-Maxwell theory with a conformally coupled and variable scalar field, in the case of cosmological constant $\Lambda>0$ for different values of the parameter $\epsilon=\Lambda/\alpha$, which determines the squared charge $Q^{2}/m^{2}$ divided by the mass $m$. The dashed line represents the unstable solutions with radius $a_{0}/m$ while there are no stable configurations. The dark gray areas have no physical meaning (see text). The shaded areas represent the regions where the WEC is satisfied.
Figure 3: Stability of a spherical thin-shell wormholes symmetric across the throat, in Einstein-Maxwell theory with a conformally coupled and variable scalar field, in case of cosmological constant $\Lambda<0$ for different values of $\epsilon=\Lambda/\alpha$, which determines the squared charge $Q^{2}/m^{2}$ divided by the mass $m$. The solid line represents static stable configurations and the dashed line shows the unstable solutions, with radius $a_{0}/m$. The shaded areas represent the regions where the WEC is satisfied.
In our construction, when $0<\Lambda m^{2}<3/16$ we require that the throat radius should satisfy $r_{h}<a_{0}<r_{c}$, to avoid the presence of the event horizon, and both the Cauchy horizon and the singularity inside it; in this case, we have a traversable wormhole with a cosmological horizon at each side of the throat. If $\Lambda m^{2}>3/16$ the only horizon is $r_{i}$, with the metric function $A(r)$ positive for $r<r_{i}$ and negative for $r>r_{i}$. This suggests that $r_{i}$ would behave as a kind of cosmological horizon, which surrounds a naked singularity at the origin; it is not clear if the attribute of being a Cauchy horizon is maintained or not in this case. We adopt $0<a_{0}<r_{i}$ to remove the singularity and also to have a positive metric function. When $\Lambda<0$ we only require that $a_{0}>0$ so the singularity at the origin is not present. We use the formalism of the previous section, so $a_{0}$ has to fulfill Eq. (31), while the stability analysis comes from Eq. (30). The scalar field shown in Eqs. (58) and (59) should be real, which imposes the restrictions (60). It is important to emphasize that the parameter $\epsilon$ does not affect the junction condition (26) nor the stability, determined by the second derivative of the potential (30). These equations involve only the metric function, which only depends on the mass and the cosmological constant. The WEC is satisfied when both inequalities in Eq. (35) hold. The parameter $\epsilon$ determines the WEC because it involves the scalar field in which the parameter $\alpha$ appears.
We graphically show the results for the case $\Lambda>0$ in Fig. 2. We analyze the existence and the stability of the static configurations, as well as the WEC, by plotting $a_{0}/m$ against $\Lambda m^{2}$ for different values of the parameter $\epsilon$ and consequently of $Q/m$. The black curve corresponds to the solutions that satisfy (31), where the stable ones are drawn with a solid line and the unstable ones with a dashed line. The dark gray areas represent regions without physical meaning while the dashed areas correspond to regions in which the WEC is satisfied. We see that there is one solution, which is unstable for the whole range of the parameters; it has $a_{0}/m>1$ when $0<\Lambda m^{2}<3/16$ and $a_{0}/m<1$ when $\Lambda m^{2}>3/16$. Additionally, the WEC is always violated at the throat if $0<\Lambda m^{2}<3/16$. However, for $\Lambda m^{2}>3/16$ when the absolute value $\epsilon$ is large enough (i. e. small enough $Q^{2}/m^{2}$) the solution satisfies WEC, but at the expense of $f(\phi_{0})<0$.
We continue now with the special case of $\Lambda=0$. The metric takes the form of the extremal Reissner-Nordström spacetime in which $Q=m$; it has a horizon at $r_{h}=m$ and the scalar field vanishes. In this case, there are no real and positive solutions of Eq. (31), then our construction is not possible.
Finally, we present the results for $\Lambda<0$ in Fig. 3, which shows the more representative results. In this case, for $(-107-51\sqrt{17})/512\approx-0.62\leq\Lambda m^{2}<0$ there is a pair of solutions of Eq. (31), the larger one is stable and the other is unstable. Depending on the value of $\epsilon$ and the other parameters these solutions can satisfy the WEC or not. For a small absolute value $|\epsilon|$ (i.e. large $Q^{2}/m^{2}$), the WEC is fulfilled in a short part of the stable curve which grows as $|\epsilon|$ increases to finally satisfy it along the complete line. After that, the unstable solution also begins to satisfy the WEC and for large enough $|\epsilon|$ it always does. Again, the throat can have normal matter but at the cost of having that $f(\phi_{0})<0$.
It is worthy to mention that the WEC is always satisfied in the bulk. The energy-momentum tensor in an orthonormal frame reads
$T^{\hat{\mu}\hat{\nu}}=\frac{m^{2}}{8\pi r^{4}}\mathrm{diag}\left(1,-1,1,1\right),$ (68)
which for a perfect fluid takes the form $T^{\hat{\mu}\hat{\nu}}=\mathrm{diag}(\rho,p_{1},p_{2},p_{3})$, with $\rho$ the volume energy density and $p_{i}$ the pressures; it is easy to check that $\rho\geq 0$ and $\rho+p_{i}\geq 0$, so the WEC is fulfilled.
To sum up, when $\Lambda>0$ we have found one solution which is always unstable; if $\Lambda=0$ our construction is not possible; while for $\Lambda<0$ we have obtained a pair of solutions, one stable and the other unstable. The configurations can satisfy the WEC or not, depending on the particular values of $\epsilon=\Lambda/\alpha$ and $\Lambda m^{2}$, but the presence of normal matter requires $f(\phi_{0})<0$. The consequence, as mentioned above, is a negative effective Newton constant and a graviton with a ghost behavior. The WEC is satisfied at the bulk in all cases.
5 Conclusions
In this work, we have constructed a class of spherical thin-shell wormholes in scalar-tensor theories with a self-interaction potential non-minimally coupled to gravity, in which the throat is located at the shell. For wormholes symmetric across the throat, we have studied the stability of static configurations under perturbations preserving the spherical symmetry, establishing a condition in terms of the second derivative of a potential.
We have found that the symmetry across the spherical throat implies that $2f^{\prime}(\phi)[K]=n^{\mu}[\partial_{\mu}\phi]=0$ and consequently $[K]=0$, which brought the equations to a similar form as in the $F(R)$ gravity counterpart. Thin-shells wormholes in $F(R)$ theories have been studied extensively in [56], in which the condition $[K]=0$ also holds [64]. The expressions for the energy density $\sigma$, the pressure $p$, and the condition that the throat radius $a$ should satisfy, have a similar form to the corresponding ones in $F(R)$ theories. They differ by a factor proportional to $f(\phi)$ instead of a function of the Ricci scalar $R$. This behavior is not surprising, since $F(R)$ gravity and scalar-tensor theories are closely related. However, the solutions to use in the wormhole construction depend on the theory adopted, so the specific models present different characteristics.
In particular, we have considered two examples of charged thin-shell wormholes with spherical symmetry, by using known black hole solutions in Einstein-Maxwell theory with a conformally self-interacting scalar field $\phi(r)$. We have studied the stability of the static configurations under radial perturbations and analyzed the fulfillment of the WEC at the throat and in the bulk. In the first example, where the Lagrangian of the theory is quadratic in the scalar field, the adopted solution has a constant $\phi$ and a metric depending on the mass $m$, the charge $Q$, and the scalar hair $s$. We have concluded that there is a pair of stable and unstable configurations for a range of values of $s/m^{2}$ and $Q/m$. The stable solutions appear when $0<s/m^{2}<1$ for large $Q>Q_{c}$ and when $1\leq s/m^{2}<9/8$ for comparatively small values of $Q/m$; in both cases with $a_{0}/m$ close to $1$. In the second example, the Lagrangian includes a cosmological constant $\Lambda$ and a term of fourth order in the scalar field, with a coupling parameter $\alpha$. The chosen solution has a variable $\phi(r)$ and a metric depending on $m$ and $\Lambda$; the charge per mass $Q/m$ is determined by $\Lambda/\alpha$. For $\Lambda>0$ we have shown that there exist one unstable solution and when $\Lambda<0$ we have found two solutions, the one with larger $a_{0}/m$ is stable while the other is unstable.
In both examples, we have verified that the existence of exotic matter cannot be avoided at the throat when requiring the physically sound condition that $f(\phi)>0$, since for a wormhole symmetric across the throat the energy conditions are always violated in any scalar-tensor theory, as it has been already proved [63].
Acknowledgments
EFE and GFA acknowledge the support by CONICET. VK thanks IAFE for the hospitality and acknowledges the financial support from Erasmus+ EU funds and the national co-financing grant from Estonian Ministry of Education and Science for Erasmus+ mobility activity.